A senior MIS design class project team has developed the following schedule of activities for their project, using their best estimate of completion times. Both written and oral reports are required. Please construct a project network to answer the following questions.

Activity

Time (days)

Immediate Predecessor

A

Find client

4

---

B

Write prospectus

2

A

C

Obtain approval from client and professor

3

B

D

Complete programming

12

C

E

Do industry background research

10

---

F

Write final paper

7

D, E

G

Write oral report

5

D, E

How many days of slack does Activity G have?

A. 10 days

B. 0 days

C. 11 days

D. 2 days.
Answer #1

Answer: D. 2 days

Explanation:

step 1 find early start, and early finish from th start of the project for each activity
Activity Duration, D Early start, ES=Max of early finish of preceeding activities Early finish, EF = ES + D
A 4 0 4
B 2 4 6
C 3 6 9
D 12 9 21
E 10 0 10
F 7 21 28
G 5 21 26
step 2 Find Maximum EF and keep it as the LF of last activities

28 is the Max EF.

step 3 calculate LS for each activity after calculating its EF's.
Activity Duration, D Early start, ES=Max of early finish of preceeding activities Early finish, EF = ES + D Late finish, LF= Min of LS of successor activities Late start, LS= LF - D Total slack= LF-EF
A 4 0 4 4 0 0
B 2 4 6 6 4 0
C 3 6 9 9 6 0
D 12 9 21 21 9 0
E 10 0 10 21 11 11
F 7 21 28 28 21 0
G 5 21 26 28 23 2

Hence activity G has slack of 2 days.